OpenStudy (anonymous):

Use the discriminant to determine which of these equations has no solution. A. x^2-x-3 = 0 B. x^2 + 10x + 25 = 0 C. x^2 + x + 1 D. x^2 -1 = 0

OpenStudy (inkyvoyd):

The discriminant D in the equation \(ax^2+bx+c=0\) is \(\Huge D=b^2-4ac\)

Parth (parthkohli):

One of them isn't even an equation. Now can you see it or you've written the question wrong?

OpenStudy (anonymous):

discriminant is : \[D=b ^{2}-4ac\] if this value is negative, means no solution

OpenStudy (anonymous):

wrote it wrong. C should be x^2 + x + 1 = 0

OpenStudy (anonymous):

i know how to do the discriminant....but I got two negatives C. and D. I got this: A. 13 B. 0 C. -3 D. -4 so what did I do?

OpenStudy (anonymous):

A. 1-4(-3)=13, correct B. 0, correct C.correct D. 0-4(-1)=4 so c is the only one with no real solutions

OpenStudy (anonymous):

wow, thanks!